# Complex Valued Deep Operator Network (DeepONet) $[\mathcal{G}]$ for Three Dimensional Maxwell’s Equations: $\mathcal{G}\in\mathbb{C}^{m\times n}$

### v1

@@ 5 Summary
In this study, we have developed a novel complex-valued DeepONet to address the challenge of operator learning for solutions expressed by complex numbers. This approach preserves the phase information of complex numbers, which offers a significant advantage over real-valued models for problems governed by inherently complex-valued equations. This study leads to important contributions that enhance the efficiency and applicability of neural operators for solving direct scattering problems governed by the time-harmonic Maxwell’s equations.
This works shows detailed numerical examples validating the effectiveness of the complex-valued DeepONet for predicting the total fields from inputs of the incident fields and the angular frequency. The proposed model beats the vanilla real-valued DeepONet in terms of prediction accuracy and maintains its performance when scaling to larger datasets. Additionally, the use of acquisition function to sample the frequency input space enables the model to perform well with only very limited training samples available.
This study not only extends the DeepONet for complex-valued data but also introduces a technique to combine multiple DeepONet models to impose the boundary conditions for PEC scatterers. We have additionally adapted the the reparametrization and two-step training method of real-valued DeepONet to the complex valued DeepONet to enhance the efficiency of model training and accuracy.
@@ Acknowledgments
We gratefully acknowledge the financial support from Hypercomp Inc. This research was partially conducted using computational resources and services at the Center for Computation and Visualization (OSCAR) at Brown University. Additionally, we thank the AFOSR Grant FA9550-23-1-0671 for supporting the purchase of modern GPU hardware.
@@
The data and code related to the work will be made available through Github repository upon acceptance of the manuscript.

### v2

@@ 5 Summary
In this study, we have developed a novel complex-valued DeepONet to address the challenge of operator learning for solutions expressed by complex numbers. This approach preserves the phase information of complex numbers, which offers a significant advantage over real-valued models for problems governed by inherently complex-valued equations. This study leads to important contributions that enhance the efficiency and applicability of neural operators for solving direct scattering problems and computing the radar cross section governed by the time-harmonic Maxwell’s equations.
This work presents detailed numerical examples demonstrating the effectiveness of the complex-valued DeepONet in predicting the total electromagnetic fields from inputs consisting of the incident fields and the angular frequency. We consider two three-dimensional geometries: a metallic sphere with a smooth surface and a more complex almond-shaped domain featuring sharp corners. In both cases, the proposed complex-valued model outperforms the conventional real-valued DeepONet in terms of prediction accuracy and maintains robust performance when scaled to larger datasets. Furthermore, by incorporating an acquisition function to strategically sample the frequency input space, the model is able to achieve high accuracy even with a very limited number of training samples, highlighting its efficiency and potential for real-time applications.
This study not only extends the DeepONet for complex-valued data but also introduces a technique to combine multiple DeepONet models to impose the boundary conditions for PEC scatterers. We have additionally adapted the reparametrization and two-step training method of real-valued DeepONet to the complex valued DeepONet to enhance the efficiency of model training and accuracy.
@@ Acknowledgments
We gratefully acknowledge the financial support from Hypercomp Inc. This research was partially conducted using computational resources and services at the Center for Computation and Visualization (OSCAR) at Brown University. Additionally, we thank the AFOSR Grant FA9550-23-1-0671 for supporting the purchase of modern GPU hardware. Finally, we thank the valuable insights and support from Prof. George Em Karniadakis at Brown University.
@@
The data and code related to the work will be made available at the GitHub repository https://github.com/Qile-J/CV_DeepONet.git.
Appendix A Comparison between Complex-Valued and Real-Valued Neural Networks
To illustrate why the complex-valued neural networks (CVNNs) are better than real-valued neural networks for learning inherently complex-valued functions, we present a simple one-layer network comparison. We consider the function $f(z)=|z|e^{ik\arg(z)}$ with $k=5$ and $z\in\mathbb{C}$.
A.1 Network architecture
The complex-valued neural network used in this example is a standard one-hidden-layer network, where all operations exist in the complex domain $\mathbb{C}$. The NN-approximated function $f_{\text{complex}}(z)$ is given as:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{\text{complex}}(z)=\mathbf{w}^{(2)}\cdot g_{\mathbb{C}}(\mathbf{W}^{(1)}z+\mathbf{b}^{(1)})+b^{(2)}}$ (26)
where the input is $z\in\mathbb{C}$, the hidden layer has $\mathbf{W}^{(1)}\in\mathbb{C}^{H\times 1}$ (complex weights with $H$ hidden units), $\mathbf{b}^{(1)}\in\mathbb{C}^{H}$ (complex biases), and $g_{\mathbb{C}}:\mathbb{C}\to\mathbb{C}$ is a complex activation function as defined in Equation 11. The output layer has $\mathbf{w}^{(2)}\in\mathbb{C}^{1\times H}$ and $b^{(2)}\in\mathbb{C}$.
The real-valued approach uses two completely separate networks that do not share any weights or information. The final function $f_{\text{real}}(z)$ combines two independent outputs for the real part $u$ and imaginary part $v$:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}f_{\text{real}}(z)=u(x,y)+i\cdot v(x,y)}$ (27)
where the input is $\mathbf{x}=\begin{pmatrix}x\\ y\end{pmatrix}\in\mathbb{R}^{2}$ (the real and imaginary parts of $z$ treated as separate features). The real part network is:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u(x,y)=\mathbf{w}_{u}^{(2)}\cdot g_{\mathbb{R}}(\mathbf{W}_{u}^{(1)}\mathbf{x}+\mathbf{b}_{u}^{(1)})+b_{u}^{(2)}}$ (28)
and the imaginary part network is:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v(x,y)=\mathbf{w}_{v}^{(2)}\cdot g_{\mathbb{R}}(\mathbf{W}_{v}^{(1)}\mathbf{x}+\mathbf{b}_{v}^{(1)})+b_{v}^{(2)}}$ (29)
where all weights and biases are real-valued ($\in\mathbb{R}$), and $g_{\mathbb{R}}$ is ReLU.
A.2 Why the Complex Network Performs Better
The function $f(z)=|z|e^{i5\arg(z)}$ is inherently polar. It takes $z=re^{i\theta}$ and maps it to $re^{i(5\theta)}$, rotating the phase by 5x while keeping the magnitude. The core operation in a complex dense layer, $W\cdot z$, naturally performs rotation and scaling:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W\cdot z=(|W|\cdot|z|)\cdot e^{i(\arg(W)+\arg(z))}}$ (30)
The network’s basic operation adds phases (rotates) and multiplies magnitudes. This aligns well with the structure of the target function.
In contrast, the real-valued networks do not operate on angles or magnitudes directly. They process $x$ and $y$ coordinates separately. The real part network must learn:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}u(x,y)=\sqrt{x^{2}+y^{2}}\cdot\cos(5\cdot\text{atan2}(y,x))}$ (31)
and the imaginary part network must learn:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}v(x,y)=\sqrt{x^{2}+y^{2}}\cdot\sin(5\cdot\text{atan2}(y,x))}$ (32)
These are complex, periodic, trigonometric functions. Approximating them with piecewise-linear ReLU activations is difficult. Furthermore, the two networks cannot share information. Each must independently learn $\sqrt{x^{2}+y^{2}}$ and $\text{atan2}(y,x)$ without coordination.
A.3 Experimental Results
We generate 10,000 training samples and 2,000 test samples by randomly sampling complex numbers from [1,1]+i[1,1][-1,1]+i[-1,1]. Both models use only one hidden layer with 64 neurons and are trained for 5,000 epochs with the Adam optimizer at learning rate $10^{-3}$. The training and testing loss histories are shown in Figure 16. The complex-valued network achieves a final test error of 4.16%, while the real-valued approach achieves 5.34%.
Figure 16: Training and testing loss histories for the complex-valued neural network (blue) and real-valued neural networks (red). The left plot shows the mean squared error loss, while the right plot shows the relative L2 error. The complex-valued network converges to a lower error.
Figure 17 presents predictions on a uniform grid. The top row shows real parts, the middle row shows imaginary parts, and the bottom row shows absolute pointwise errors. The complex-valued network (mean error: 0.0289) achieves lower errors than the real-valued approach (mean error: 0.0352).
Figure 17: Comparison of predictions for $f(z)=|z|e^{i5\arg(z)}$. The complex-valued network captures the intricate patterns around the origin more accurately than the real-valued approach.
This example suggests that for problems with inherent phase and magnitude structure, complex-valued neural networks may offer advantages. This motivates the design of complex-valued DeepONets to solve PDEs with complex-valued solutions.